OpenStudy (giss10):

Find the percent of decrease. Round to the nearest tenth, if necessary. From 49 ft to 22 ft

OpenStudy (stefrheart):

First you need to subtract the original numbers: \[49-2\] Then you will take the difference of the two numbers and divide it by the original answer (49) \[49\div x\] The answer of that will then be multiplied by 100 and round it by the nearest tenth

OpenStudy (stefrheart):

*22 I made a mistake on my first equation, I put 2 and it should be 22

OpenStudy (stefrheart):

Did that make any sense?

OpenStudy (giss10):

Yes thank you!

OpenStudy (sshayer):

\[\frac{ 49-22 }{ 49 }~\times ~100=?\]

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)